# Procountable groups are not classifiable by countable structures

### v1

@@ Abstract.
We prove that the relation of topological isomorphism on procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ that expresses that two sequences of reals have bounded difference is Borel reducible to it. This marks progress on an open problem of [15], to determine the exact complexity of isomorphism among all non-archimedean Polish groups.
S. Gao acknowledges the partial support of their research by the Fundamental Research Funds for the Central Universities and by the National Natural Science Foundation of China (NSFC) grant 12271263. Nies acknowledges support by the Marsden Fund of New Zealand, UOA-1931. Both S. Gao and A. Nies attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this work was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center. G. Paolini was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA and by INdAM Project 2024 (Consolidator grant) “Groups, Crystals and Classifications”.
@@ 1. Introduction
Suppose we are given a concrete presentation of abstract mathematical objects of a certain type. It is of interest to know the complexity of determining whether two presentations describe the same abstract object. For instance, finite graphs can be concretely described by binary strings that encode adjacency matrices. The graph isomorphism problem asks whether two such strings describe the same graph. This problem is in NP, and its exact complexity has been the object of much recent study [1, 30]. The triviality problem for groups is to determine whether a finite presentation of a group decribes the trivial group. This is undecidable [29]. Therefore the $\Sigma^{0}_{1}$ relation of isomorphism between finitely presented groups is undecidable. In fact every $\Sigma^{0}_{1}$ equivalence relation is computably reducible to it ([22]; also see [25, Cor. 2.6]).
The presentations in the previous two examples are finite. However, for many classes of mathematical objects, a countably infinite amount of information is required to describe a member. In this case, one can usually view the descriptions as points in a so-called standard Borel space (such as the unit interval). Borel reducibility, introduced in [7], is the standard tool for comparing the complexity of equivalence relations on standard Borel spaces. For instance, countable graphs can be concretely presented as symmetric, irreflexive relations on $\mathbb{N}$, and their isomorphism is a well-known benchmark equivalence relation [8] denoted by GI. The isomorphism relation for a class of concretely presented objects is reducible to GI if and only there is a Borel assignment of countable structures as complete invariants to presentations of the objects in question.
The programme of [15]. A Polish group is a topological group whose topology is Polish, i.e., separable and completely metrizable. A topological group is called non-archimedean if it has a neighbourhood basis of the identity consisting of open subgroups [2]. Such groups are totally disconnected. Let $\mathrm{Sym}(\mathbb{N})$ denote the group of all permutations of $\mathbb{N}$ equipped with the pointwise convergence topology. Then $\mathrm{Sym}(\mathbb{N})$ is a non-archimedean Polish group. Up to topological isomorphism, the non-archimedean Polish groups are precisely the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ [2]. It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space (see [15]). So we can take the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ as the concrete presentations of non-archimedean Polish groups. In Kechris et al. [15] the following programme was initiated:
@@
There has been some progress on this program. We summarize the results that are known to us below.
  • One class considered in [15] was the compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bireducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel.
  • One can also consider the subclass of abelian profinite groups. Due to the Pontryagin duality, the isomorphism relation of abelian profinite groups is Borel bireducible with the isomorphism of countable torsion abelian groups. It is well known that the latter relation is classified by the Ulm invariants, and is known to be a complete analytic set and hence non-Borel (see [7, (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$.
  • The isomorphism relation of oligomorphic groups was shown to be essentially countable in [23]; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see [24, 26]; in particular in [26] it was shown that under the assumption of weak elimination of imaginaries the relation is smooth.
  • The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in [4]. Work in progress of Ruiwen Li would show that this relation is Borel bireducible with GI.
Our work marks considerable progress on an open problem of Kechris et al. [15] pertaining to this programme. In [15], the authors gave a criterion determining when isomorphism on an invariant Borel class of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ can be classified by countable structures. They asked at the end of their introduction for the complexity of isomorphism for the class of all non-archimedean Polish groups, observing that this relation is analytic and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of procountable groups already has a complexity much higher than GI.
@@
We note that the programme of [15] also asked to first determine whether a class $\mathcal{C}$ of interest is indeed Borel. While this is often straightforward from the definition, in some cases it takes work to establish. The Borelness is unknown for the class of topologically finitely generated groups, and known to be false for some classes of interest. For instance, the class of all CLI groups (those admitting a compatible complete left-invariant metric) was shown to be coanalytic and non-Borel in [18], and in [28] it is shown that four properties of closed abelian subgroups of $\mathrm{Sym}(\mathbb{N})$ are complete coanalytic and therefore non-Borel: separability (in the sense of abelian group theory), torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
The main result.
@@
A topological group $G$ is called procountable if $G$ is topologically isomorphic to the inverse limit of an inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ of countable groups and onto homomorphisms $p_{n}\colon G_{n+1}\to G_{n}$, where the topology of $G$ is the subspace topology given by the product $\prod_{n}G_{n}$, where each $G_{n}$ is equipped with the discrete topology.
It is easily seen that each procountable group is non-archimedean Polish, thus it is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. We will verify in 2.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of $1$ of normal open subgroups (see, e.g., [12, Theorem 3.3] or [19, Lemma 2]). Since each countable discrete group is procountable, we know that GI is Borel reducible to the isomorphism of procountable groups through any of the following results about countable groups:
  • The isomorphism of countable locally finite groups that are nilpotent of class 22 is Borel bireducible with GI (Mekler [21], who proves this in fact for groups of exponent 3; also see [7, (2.3)]);
  • The isomorphism of countable torsion-free abelian groups is Borel bireducible with GI ([27]);
  • The isomorphism of countable omnigenous locally finite groups is Borel bireducible with GI ([10]).
In particular, we get that the isomorphism relation of procountable groups is a complete analytic set and hence non-Borel. In this paper we address the problem of its complexity in the Borel reducibility hierarchy.
Recall that $E_{1}$ is the Borel equivalence relation of eventual equality between sequences of real numbers. By a celebrated theorem of Kechris and Louveau [14], $E_{1}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. It follows that any equivalence relation which Borel reduces $E_{1}$ has the same property. Recall that $\ell_{\infty}$ is the Borel equivalence relation on $\mathbb{R}^{\mathbb{N}}$ given by
@@
By a theorem of Rosendal (see e.g. [8, Theorem 8.4.2]), $E_{1}$ is Borel reducible to $\ell_{\infty}$, and thus $\ell_{\infty}$ is not Borel reducible to the orbit equivalence relation of any Borel action of a Polish group. The following is the main result of this paper.
@@
The equivalence relation $\ell_{\infty}$ is Borel reducible to the topological isomorphism relation on procountable groups. Consequently, the topological isomorphism relation on procountable groups is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, it is not Borel reducible to GI.
@@
    Section 4 reduces this uniform homeomorphism relation to the topological isomorphism relation on the space of procountable groups.
The second step only uses a very particular type of groups that could be called “solenoid groups”. Any group $G_{n}$ in the inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ is isomorphic to a free Coxeter group generated by an infinite set $V_{n}$, namely $\langle v\in V_{n}\mid v^{2}=e\rangle$. The binding maps are given by surjections $V_{n+1}\to V_{n}$.
@@
We close the introduction with the following open problems. The first two will be discussed again in see Section 3.4.
@@
Is the relation of topological isomorphism on procountable groups a universal analytic equivalence relation (i.e., any analytic equivalence relation is Borel reducible to it)?
@@
Is the relation of uniform homeomorphism on ultrametric Polish spaces a universal analytic equivalence relation?
@@
Is there a nilpotent variety of groups $\mathbf{V}$ such that the relation of topological isomorphism on procountable groups in $\mathbf{V}$ is not classifiable by countable structures? In particular, are abelian procountable groups classifiable by countable structures?
It follows from [27] that GI is a lower bound for topological isomorphism on procountable abelian groups.
2. Presentations of procountable groups
Recall Definition 1.1. Encoding inverse systems by countable structures with infinitely many sorts, we obtain a standard Borel space $\mathcal{P}$ of procountable groups. We show in this section that these presentations can be converted in a Borel way into procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$, and that the latter form a Borel set $\mathcal{C}$.
Notation 2.1.
$s]=\left\{x\in\mathrm{Sym}(\mathbb{N})\colon\,s\subseteq x\right\}.$
[…]
Remark 2.2.
Proposition 2.3.
Proof.
In the following, letters $s,t$ etc. range over sequences that have an extension in $G$. Recalling 2.1, we claim that (1) is equivalent to the Borel condition
Remark 2.4.
[…]
@@ 3.1. Some facts on uniform homeomorphisms
The following is a direct consequence of the definition of uniform continuity.
@@
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$ be continuous. Then $\Phi$ is uniformly continuous if and only if for all $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that for any $x,y\in[T]$, if $x\!\!\upharpoonright\!m=y\!\!\upharpoonright\!m$, then $\Phi(x)\!\!\upharpoonright\!n=\Phi(y)\!\!\upharpoonright\!n$.
@@ 3.4. Further remarks
Here we remark that our result does not show that $\cong_{u}$ is universal. By 3.3 GI is reducible to $\cong_{u}$, and by the main result 1.2 $\ell_{\infty}$ is reducible to $\cong_{u}$. However, the equivalence relation $c_{0}$ does not reduce to $\mbox{\rm GI}\times\ell_{\infty}$. Recall that $c_{0}$ is the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined as
@@
Toward a contradiction, assume there was a map $f$ witnessing that $c_{0}$ is Borel reducible to $\mbox{\rm GI}\times\ell_{\infty}$. The map $f$ induces a homomorphism of equivalence relations from $c_{0}$ to GI. By the Hjorth turbulence theorem (see [13, Theorem 3.18 and Example 3.23]; also see [8, Theorems 10.4.2 and 10.5.2]), there is a $c_{0}$-invariant comeager set $C$ on which $c_{0}$ is sent by $f$ to a single GI-class. This implies that $c_{0}$ on $C$ is Borel reducible to $\ell_{\infty}$. However, $c_{0}\upharpoonright C^{2}$ is ${\bf\Pi}^{0}_{3}$-complete (see [8, Exercise 8.5.1]) and as such, it is not Borel reducible to any $F_{\sigma}$ equivalence relation (see [8, Theorem 8.5.2] and its proof).
@@
4. Borel reduction of uniform homeomorphism to isomorphism of procountable groups
In this section we implement the second step of our proof strategy toward a proof of our Main Result 1.2, namely we exhibit a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of procountable groups.
For a countably infinite set $V$, we define the free product of cyclic groups of size $2$ with generators in $V$ by
$L(V)=\langle s\mid s^{2}=e\ \ (s\in V)\rangle.$
(This group is also known as the free Coxeter group of rank $\aleph_{0}$.) By the general theory of free products of group (see, e.g., [17, IV.4.7]), each involution is conjugate to a generator, and no two generators are conjugate. If $r\colon V\to W$, then we have a map $\widehat{r}\colon L(V)\to L(W)$ induced by $v\mapsto r(v)$.
Given a pruned tree $T$, let $T_{n}=\{s\in T\colon|s|=n\}$ be the $n$-th level of the tree $T$, and let $r_{n}\colon T_{n+1}\to T_{n}$ be the predecessor map. We may assume that $T_{1}$ is infinite because the trees we obtain through the coding in 3.4 have that property. Let

### v2

@@ Abstract.
We prove that topological isomorphism on $2$-nilpotent procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ that expresses that two sequences of reals have bounded difference is Borel reducible to it. This marks substantial progress on an open problem of [16]: to determine the exact complexity of the isomorphism relation among all non-archimedean Polish groups.
S. Gao acknowledges the partial support of their research by the Fundamental Research Funds for the Central Universities and by the National Natural Science Foundation of China (NSFC) grant 12271263. A. Nies acknowledges support by the Marsden Fund of New Zealand, UOA-1931. Both S. Gao and A. Nies attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this research was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center. G. Paolini was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA and by INdAM Project 2024 (consolidator grant) “Groups, Crystals and Classifications”.
@@ 1. Introduction
It is often useful to have concrete presentations of abstract mathematical objects of a certain type. In this case one is interested in the complexity of whether two presentations present the same abstract object. For instance, finite graphs can be concretely presented by adjacency matrices (written as binary strings). The graph isomorphism problem asks whether two such matrices present the same graph. This problem is in NP, and its exact complexity has been the object of much recent study [1, 31]. The triviality problem for groups is to determine whether a presentation given by finitely many generators and relators presents the trivial group. This is undecidable [30], so the $\Sigma^{0}_{1}$ relation of isomorphism between finite presentations of groups is undecidable. In fact, every $\Sigma^{0}_{1}$ equivalence relation is computably reducible to it by [23] (also see [26, Cor. 2.6]).
For many types of mathematical objects, a countably infinite amount of information is required to present a member. In this case, one can usually view the presentations as points in a so-called standard Borel space (such as the unit interval). Borel reducibility, introduced in [7], is the standard tool for comparing the complexity of equivalence relations on standard Borel spaces. For instance, countable graphs can be concretely presented as symmetric, irreflexive relations on $\mathbb{N}$, and their isomorphism is a well-known benchmark equivalence relation [8] denoted by GI. The isomorphism relation for a class of concretely presented objects is reducible to GI if and only there is a Borel assignment of countable structures as complete invariants to presentations of the objects in question.
A programme of Kechris et al. A Polish group is a topological group with a Polish topology, namely, one that is separable and completely metrizable. A topological group is called non-archimedean if it has a neighbourhood basis of the identity consisting of open subgroups [2]. Such groups are totally disconnected. Let $\mathrm{Sym}(\mathbb{N})$ denote the group of all permutations of $\mathbb{N}$ equipped with the pointwise convergence topology. Then $\mathrm{Sym}(\mathbb{N})$ is a non-archimedean Polish group. Up to topological isomorphism, the non-archimedean Polish groups are precisely the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ [2]. It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space (see [16]). So we can take the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ as the concrete presentations of non-archimedean Polish groups.
It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space; see for instance Kechris et al. [16]. In that article they initiated the following programme:
@@
There has been some progress on this program. We summarize the results known to us. One class considered in [16] is the class of compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bi-reducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel. One can also consider the subclass of abelian profinite groups. Due to Pontryagin duality, the isomorphism relation between abelian profinite groups is Borel bi-reducible with the isomorphism relation between countable torsion abelian groups. The latter relation is classified by Ulm invariants; it is also a complete analytic set and hence non-Borel (see [7, (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$. The isomorphism relation of oligomorphic groups was shown to be essentially countable in [24]; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see [25, 27]; in particular in [27] it was shown that under the assumption of weak elimination of imaginaries the relation is smooth. The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in [4], where it is shown that the setwise equality of infinite sequences of reals is Borel reducible to it. Work in progress of Ruiwen Li, in conjunction with [4], would show that in fact GI is reducible to it.
@@
We note that the programme of [16] asked to first determine whether a given class of interest $\mathcal{C}$ of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ is indeed Borel. While this is often straightforward from the definition, in some cases it takes work to establish. Borelness is unknown for the class of topologically finitely generated groups, and known to fail for some classes of interest. For instance, the class of all CLI groups (those admitting a compatible complete left-invariant metric) was shown to be properly co-analytic by Malicki [19]. Paolini and Shelah [29] proved that four properties of closed abelian subgroups of $\mathrm{Sym}(\mathbb{N})$ are complete co-analytic and therefore non-Borel: separability (in the sense of abelian group theory), torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
Countable groups are up to isomorphism the discrete closed subgroups $G$ of $\mathrm{Sym}(\mathbb{N})$, and being discrete is Borel because it means that the identity permutation is isolated in $G$. If a class of discrete subgroups of $\mathrm{Sym}(\mathbb{N})$ is Borel, its isomorphism relation is Borel reducible to GI. The question is whether conversely, GI reduces to the isomorphism relation on the class. Mekler [22] (as noted in [7, 2.3]) established this for the countable groups that are nilpotent of class $2$, and of exponent a fixed odd prime. His method of coding graphs into such groups will be instrumental for our main result, Theorem 1.2. Paolini and Shelah [28] showed it for countable torsion-free abelian groups. Most recently, Gao and Li [10] did so for countable omnigenous locally finite groups (an uncountable class of groups, each generalising Hall’s universal locally finite group).
The main result. Our work marks considerable progress on an open problem of Kechris et al. [16] pertaining to their programme. They asked at the end of their introduction for the complexity of isomorphism on the class of all non-archimedean Polish groups, observing that this relation is analytic, and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of 2-nilpotent procountable groups of exponent a fixed odd prime $p$ already has a complexity much higher than GI.
@@
A topological group $G$ is called procountable if $G$ is topologically isomorphic to the inverse limit of an inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ of countable groups and onto homomorphisms $p_{n}\colon G_{n+1}\to G_{n}$. The topology of $G$ is the subspace topology given by the product $\prod_{n}G_{n}$, where each $G_{n}$ is equipped with the discrete topology.
It is easily seen that a group of the form $\prod_{n}G_{n}$, and hence each procountable group, is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. For background and further equivalent characterisations, see e.g., [20]. We will verify in 3.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of the neutral element consisting of normal open subgroups (see, e.g., [12, Theorem 3.3] or [20, Lemma 2]). The procountable groups satisfying some law (such as being 2-nilpotent) then also form a Borel set: from a closed subgroup $G$ of $\mathrm{Sym}(\mathbb{N})$, using [14, Th. 12.13] one can in a Borel way determine a countable dense subgroup $D$, and it satisfies to require that $D$ satisfies the law.
Since each countable discrete group is procountable, by the aforementioned result of Mekler [22], GI is reducible to the isomorphism relation on 2-nilpotent procountable groups of exponent $p$. It follows that the topological isomorphism relation of procountable groups is a complete analytic set and hence non-Borel.
Recall that $E_{1}$ is the Borel equivalence relation of eventual equality of sequences of real numbers. By a celebrated theorem of Kechris and Louveau [15], $E_{1}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. It follows that any equivalence relation to which $E_{1}$ is Borel reducible has the same property. Recall that $\ell_{\infty}$ is the Borel equivalence relation on $\mathbb{R}^{\mathbb{N}}$ given by
@@
By a result of Rosendal (see [8, Theorem 8.4.2]), $E_{1}$ is Borel reducible to $\ell_{\infty}$, and thus $\ell_{\infty}$ is not Borel reducible to the orbit equivalence relation of any Borel action of a Polish group. The following is the main result of this paper.
@@
Fix an odd prime $p$. The equivalence relation $\ell_{\infty}$ is Borel reducible to the topological isomorphism relation on 2-nilpotent procountable groups of exponent $p$. Consequently, the topological isomorphism relation on such procountable groups is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, it is not Borel reducible to GI.
@@
    Section 4 reduces this uniform homeomorphism relation to the topological isomorphism relation on the Borel space of procountable nilpotent groups of class $2$ and exponent $p$, for $p$ an odd prime.
The second step only uses a very particular type of groups that could be called “solenoid groups”. All the groups $G_{n}$ in the inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ are isomorphic. Each group $G_{n}$ is generated by a set $V_{n}$, and the binding maps $p_{n}$ are determined by surjections $V_{n+1}\to V_{n}$.
@@
We close the introduction with some open questions. The first two will be discussed again in see Section 2.4. An analytic equivalence relation on a Borel space is called universal analytic if each analytic equivalence relation is Borel reducible to it.
@@
Is the equivalence relation of topological isomorphism on procountable groups universal analytic?
We note that our original construction for the second step used as $G_{n}$ the free Coxeter group generated by $V_{n}$, namely, the group freely generated by involutions $g_{v}$ for $v\in V_{n}$; see the arxiv version 2512.12256.v1. That approach does not yield nilpotent-2 groups. However, it is somewhat simpler to verify, because the only elements of order 2 in $G_{n}$ are the conjugates of the generators, so that the elements of $V_{n}$ up to conjugacy can be recovered from $G_{n}$. It is possible that this method could lead to an affirmative answer Question 1.4.
@@
Is the equivalence relation of uniform homeomorphism on ultrametric Polish spaces universal analytic?
@@
Are the abelian procountable groups classifiable by countable structures?
Notice that it follows from [28] that GI is a lower bound for the complexity of this classification problem.
@@ 2.1. Some facts on uniform homeomorphisms
The following is a direct consequence of the definition 2.1 of the metric on path spaces, and the definition of uniform continuity.
@@
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$. Then $\Phi$ is uniformly continuous if and only if for all $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that for any $x,y\in[T]$, if $x\!\!\upharpoonright\!m=y\!\!\upharpoonright\!m$, then $\Phi(x)\!\!\upharpoonright\!n=\Phi(y)\!\!\upharpoonright\!n$.
@@ 2.4. Further remarks
We remark that our result does not show that $\cong_{u}$ is universal. By 2.3 GI is reducible to $\cong_{u}$, and by the main result 1.2 $\ell_{\infty}$ is reducible to $\cong_{u}$. However, the equivalence relation $c_{0}$ does not reduce to $\mbox{\rm GI}\times\ell_{\infty}$. Recall that $c_{0}$ is the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined as
@@
Toward a contradiction, assume that there is a map $f$ witnessing that $c_{0}$ is Borel reducible to $\mbox{\rm GI}\times\ell_{\infty}$. The map $f$ induces a homomorphism of equivalence relations from $c_{0}$ to GI. By the Hjorth turbulence theorem (see [13, Theorem 3.18 and Example 3.23]; also see [8, Theorems 10.4.2 and 10.5.2]), there is a $c_{0}$-invariant comeager set $C$ on which $c_{0}$ is sent by $f$ to a single GI-class. This implies that $c_{0}$ on $C$ is Borel reducible to $\ell_{\infty}$. However, $c_{0}\upharpoonright C^{2}$ is ${\bf\Pi}^{0}_{3}$-complete (see [8, Exercise 8.5.1]) and as such, it is not Borel reducible to any $F_{\sigma}$ equivalence relation (see [8, Theorem 8.5.2] and its proof).
@@
3. Preliminaries on procountable groups
In 1.1 we defined procountable groups. This section provides material on procountable groups that will be needed for the second step in our overall proof of the Main Result 1.2.
3.1. Presentations of procountable groups
Encoding inverse systems by structures with domain $\mathbb{N}$ that have infinitely many sorts, we obtain a standard Borel space $\mathcal{P}$ of presentations of procountable groups. We show that these presentations can be converted in a Borel way into procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$, and that the latter form a Borel set $\mathcal{C}$.
Notation 3.1.
$s]=\left\{x\in\mathrm{Sym}(\mathbb{N})\colon\,s\subseteq x\right\}.$